OpenStudy (jkbo):

Please help.What is the solution to the system of equations? x-3y-z=-9 -2x+y+2z=3 2x+y+3z=8 a.(1, 3, 1) b.(1, –3, 1) c.(–1, 3, 1) d.(1, 3, –1)

OpenStudy (primeralph):

Eliminate y from 2 and 3.

OpenStudy (texaschic101):

you need help ?

OpenStudy (primeralph):

@texaschic101 Just go ahead.

OpenStudy (texaschic101):

you think I should just do it ?

OpenStudy (primeralph):

We need your wisdom.

OpenStudy (texaschic101):

x - 3y - z = -9 -2x + y + 2z = 3 -->(3) --------------- x - 3y - z = -9 -6x + 3y + 6z = 9 (result of multiplying by 3) ---------------add -5x + 5z = 0 -2x + y + 2z = 3 -->(-1) 2x + y + 3z = 8 ---------------- 2x - y - 2z = -3 (result of multiplying by -1) 2x + y + 3z = 8 ----------------add 4x + z = 5 -5x + 5z = 0 4x + z = 5 --->(-5) ------------- -5x + 5z = 0 -20x - 5z = -25 (result of multiplying by -5) --------------add -25x = -25 x = 1 now sub 1 in for x 4x + z = 5 4(1) + z = 5 4 + z = 5 z = 5 - 4 z = 1 now sub known values into any of the original equations x - 3y - z = 9 1 - 3y - 1 = 9 -3y = 9 y = -3 check... x - 3y - z = 9 1 - 3(-3) - z = 9 1 +9 -1= 9 9 = 9 (correct) x = 1, y = 3, z = 1 or (1,3,1)

OpenStudy (primeralph):

@texaschic101 I admire your patience.

OpenStudy (texaschic101):

it took a while....I messed up....but I finally fixed it and got the answer

OpenStudy (jkbo):

thank you both